RC Meeting, Saturday, October 02 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. Beam has been mostly stable, and quality has been improving gradually.

Hall C Status

  1. No showstoppers reported
  2. I haved scheduled SHMS rotation to 5.5 deg for Sunday Oct 03 09:00 am (tomorrow)

Current Run Plan

  1. Oct 02 Day/Swing --> Oct 03 Owl/Day:
    • Continue data-taking at current PION-LT kinematic setting ( low eps, Q2=6 GeV2, W=3.19, x=0.39 ) until tomorrow Sunday 09:00 am
      • As of Oct 02 mid-Owl shift, we have accumulated ~85 % of the approved beam charge
      • Shift crew has requested beam current to be lowered to 45 uA for the remainder of this kinematic setting
    • As of End-of-Owl Shift today, we have accumulated (thanks Jacob):
      • 8669 Pion Events (86.7%), 19997.68 mC charge delivered (87.1%)
        where the total gained across the shift are: 589 Pion Events gained (5.9%), 1306.02 mC charge delivered gained (5.7%)
  2. SHMS Rotation to 5.5 deg has been scheduled for a "Controlled Access" Oct 03 Sunday Morning at 9:00 am ( CLEARED )
    • Have contacted Amy Comer/Steve Lassiter and it was concluded that rotation does require 2 techs for beamline spotting during rotation
      • 1 tech on top of HMS platform + 1 tech on floor (in-between HMS/SHMS spectrometers as closes as possible to the pivot area)
      • Garth will volunteer to be a beamline spotter in case additional beamline spotters are needed
    • Have contacted Walter Kellner and the 2 techs can ONLY come in on Sunday
    • Have contacted Dave Hamlette (RADCON), and they will provide support as well
    • Brad will also be in the Hall during this access to install spare HCCAM03 as the previous one has failed --> HCCAM03 Broken

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  2. HCCAM03 Hall Camera has failed and needs replacement --> HCCAM03 Broken ( PENDING )
    • Brad will come in during the Hall C "Controlled Access" scheduled on Sunday 9:00 am to install spare camera.

Special Requests

  1. Contact Ellen Becker if there is an opportunity to access the hall (longer than 4 hours). This is for diagnosing the SHMS power supply. ( PENDING )